OpenStudy (anonymous):

lim as x approaches 4 from the left, (x+4)/(x^2-16)

OpenStudy (anonymous):

lim as x approaches -4, (x+4)/(x^2-16) lim as x approaches 0, for the same function lim as x approaches infinity

OpenStudy (abb0t):

Take note of the denominator: \[\lim_{x \rightarrow 4^-} \frac{ x+4 }{ x^2-16 } \] \[x^2-16 = (x+4)(x-4)\] cancel like terms. That should help guide you in the right direction.

OpenStudy (anonymous):

i have 1/0 as answer, does that mean the limit does not exist ?

OpenStudy (abb0t):

It's approaching from -∞

OpenStudy (anonymous):

f(x)= 1/(x-4) thats the function i use for all the problems

OpenStudy (abb0t):

Yeah. But it's approaching 4 from the left, meaning negative. That's why it's -∞

OpenStudy (anonymous):

so for x approaching -4 i would get 1/(-4-4) which is 1/-8 how would i write the limit for that

OpenStudy (anonymous):

and x approaching 0 i would get 1/-4 how would i state the limit for that and as x approaches infinity i really don't get it